OpenStudy (anonymous):

-Calculate the limit: 1. lim x→∞ (ln(3x+1) -ln(2x+1)) 2. lim x→∞ (ln sqrt 5x^2+2)-lnx)

hartnn (hartnn):

apply the log rules you know ln A -ln B =... ?

OpenStudy (anonymous):

= a-b ?

hartnn (hartnn):

noo....ln A-ln B = ln (A/B) so, what about ln(3x+1) -ln(2x+1) ?

OpenStudy (anonymous):

3x+1 divided by 2x+1 Answer is x+1 ?? am I right?

hartnn (hartnn):

nopes, from both numerator and denominator of (3x+1)/ (2x+1) can you factor out the x ?

OpenStudy (anonymous):

if i factor x from top and bottom, i'd have x(2x+1)/ x(x+1) on the bottom

OpenStudy (anonymous):

ok im stuck :/